A portfolio consists of 280 shares of Stock C that sells for $53 and 245 shares of Stock D that sells for $22. What is the portfolio weight of Stock C?

.7947

.2284

.2664

.7336

.8253
Answer #1

Total value of C=(280*53)=$14840

Total value of D=(245*22)=$5390

Total value =$(14840+5390)=$20230

Hence portfolio weight of C=(14840/20230)

=0.7336(Approx).